Session: Feb 2005 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Hartebeesthoek (Jonathan Quick at Wed Mar 2 12:08:35 2005) Minor failures : RFI
Intermittent RFI seen in VC's 5 & 6.

Torun (K. Borkowski at Wed Mar 2 12:20:29 2005) Minor failures :
BBC3 unlocked all the time (problem not related to the BBC unit itself, but electronics prior to it). No other known problems.

Medicina (Marco Poloni at Wed Mar 2 12:33:57 2005) Success :

Westerbork (Tony Foley at Thu Mar 3 10:41:32 2005) Minor failures : WIND HIGH
Wb used in phased array mode. SNOW covering dished and one got stuck. Tsys raised as snow covered the surface. Also one telescope out with receiver oscillating

Jodrell_Bank (Alastair Gunn at Mon Mar 7 11:58:33 2005) Success :
The Lovell Telescope was used. No known problems.

Noto (P. Cassaro at Mon Mar 7 15:27:34 2005) Success :
No known problems

Effelsberg (A. Kraus at Fri Mar 11 12:29:12 2005) Minor failures :
RFI in VCs 5-8.